OpenStudy (anonymous):

Ben has 15 more music CD's than sam. Together they have a total of 55 CD's. How many CD's does Sam have?

OpenStudy (anonymous):

let Sam has x CDs Ben then has x+15 CDs as the total is 55 so x+x+15=55 2x+15=55 2x+15-15=55-15 2x=40 x=20 therefore, sam has 20 CDs

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

you are welcome :D